OpenStudy (anonymous):

ASAP! If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x + 5y – 2z = –1 x + 3y + z = –2 3x – 2y + 3z = –17 A. –y – 4z = 3 –11y = –11 B. 4x + 11y = –5 –11y = –11 C. –y = –5 6x + 7y = –11 D. 11y = 3 7y + 6z = –11 PLEASE ANSWER ASAP

OpenStudy (anonymous):

so if the second equation has to have x isolated, start with that x + 3y + z = –2 becomes x=3y+z-2 you then want to substitute this into the other 2 equations 2(3y+z-2) + 5y – 2z = –1 => 6y+2z-4+5y-2z=-1 that becomes 11y=3 the thirs equation is 3(3y+z-2) – 2y + 3z = –17 => 9y+3z-6-2y+3z=-17 that becomes 7y+6z=-11 i believe that your answer is D